Determine the electrical force of attraction between two balloons that are charged with the opposite type of charge but the same quantity of charge. The charge on the balloons is 6.0 x 10-7 C and they are separated by a distance of 0.50 m.



1
Expert's answer
2022-04-17T17:06:22-0400
"F=k\frac{q_1q_2}{r^2}\\
=9*10^9*\frac{6.0*10^{-7}*6.0*10^{-7}}{0.50^2}=0.013\:\rm N"